During high tide, a gate is held in place as shown by a stopped at D.
C4 4 m 2 m During high tide, a gate is held in place as shown by a stopper at D. Determine the resultant forces (in N/m since we do not know the length of the gate) due to water on side A and on side B of the gate. Water = 9810 N/m?
Answer #1

Pressure at the bottom on A side = rho g h = 1000*9.81*3 Pa.

Average pressure = 1000*9.81*3*0.5

Force from A = 1000*9.81*3*0.5*3 = 44145 N/m

Force from B = 1000*9.81*2*0.5*2 = 19620 N/m

Net force = 44145 - 19620 = 24525 N/m   Answer